### v2

# Leveraging Online Olympiad-Level Math Problems for LLMs Training and Contamination-Resistant Evaluation

Abstract
Advances in Large Language Models (LLMs) have sparked interest in their ability to solve Olympiad-level math problems. However, the training and evaluation of these models are constrained by the limited size and quality of available datasets, as creating large-scale data for such advanced problems requires extensive effort from human experts. In addition, current benchmarks are prone to contamination, leading to unreliable evaluations. In this paper, we present an automated pipeline that leverages the rich resources of the Art of Problem Solving (AoPS) forum, which predominantly features Olympiad-level problems and community-driven solutions. Using open-source LLMs, we develop a method to extract question-answer pairs from the forum, resulting in AoPS-Instruct, a dataset of more than 600,000 high-quality QA pairs. Our experiments demonstrate that fine-tuning LLMs on AoPS-Instruct improves their reasoning abilities across various benchmarks. Moreover, we build an automatic pipeline that introduces LiveAoPSBench, an evolving evaluation set with timestamps, derived from the latest forum data, providing a contamination-resistant benchmark for assessing LLM performance. Notably, we observe a significant decline in LLM performance over time, suggesting their success on older examples may stem from pre-training exposure rather than true reasoning ability. Our work presents a scalable approach to creating and maintaining large-scale, high-quality datasets for advanced math reasoning, offering valuable insights into the capabilities and limitations of LLMs in this domain. Our benchmark and code is available at https://livemathbench.github.io/leaderboard.html
1 Introduction
Large language models (LLMs) have shown tremendous success in solving various tasks such as code generation (Li et al., 2022), math reasoning (Shao et al., 2024), and commonsense reasoning (Zellers et al., 2019; Achiam et al., 2023), suggesting that current models may show signs of artificial general intelligence (AGI) (Bubeck et al., 2023). Math reasoning is perhaps one of the most challenging tasks for the LLMs, since mathematics is inherently structured, requiring not just recall of facts but also rigorous logical inference, abstraction, and understanding of formal symbolic systems. As such, there have been grand challenges (Selsam et al., 2019) and million-dollar prizes AIMO (2023) established for a model capable of solving Olympiad-level math problems.
On the training side, despite significant progress in certain areas, such as geometry, particularly with the assistance of symbolic methods (Trinh et al., 2024), the performance of LLMs remains limited on Olympiad-level problems (He et al., 2024). One of the key challenges in advancing competition-level math reasoning, compared to other domains like coding or grade-school math, is the scarcity of large-scale data. Creating valid and challenging math questions, along with providing correct solutions, is costly. This is especially true for Olympiad-level problems, which can be time-consuming even for experts. This highlights the need for scalable and automated methods to collect high-quality data for Olympiad-level problems to facilitate further advancements in this field.
On the evaluation side, in contrast to the rapid advancements in LLMs, the evaluation of their math reasoning capabilities remains relatively underdeveloped. First, as aforementioned, the cost of creating and annotating advanced math problems is high. Second, popular datasets such as MATH (Hendrycks et al., 2021b) and GSM8K (Cobbe et al., 2021) have been saturated by both open-source and closed-source models (Achiam et al., 2023; Yang et al., 2024b). Third, benchmarks whose test sets are publicly available online (Hendrycks et al., 2021b; Cobbe et al., 2021; He et al., 2024; Zhang et al., 2023b) are prone to potential contamination. Although techniques like n-gram matching and locality-sensitive hashing have been applied as a common practice (Achiam et al., 2023; Dubey et al., 2024; Yang et al., 2024a) to reduce contamination, they still suffer low accuracy and would not be able to rule out rephrased questions, as shown by Yang et al. (2023). Given these limitations, it is crucial to develop an evolving evaluation benchmark that contains abundant and up-to-date test samples, and designed with appropriate difficulty to fairly assess a model’s math reasoning abilities.
The Art of Problem Solving111https://artofproblemsolving.com/community (AoPS) forum is a rich resource for Olympiad-level math problems, featuring discussions on topics such as algebra, geometry, combinatorics, and number theory from competitions like AMC (AOPS, 2023), AIME (AOPS, 2024), and the International Mathematical Olympiad (IMO). However, the forum’s unstructured nature, including irrelevant comments and incomplete solutions, poses challenges in extracting high-quality, structured question-answer (QA) pairs. Developing an effective automated pipeline to curate these QA pairs is essential to address the scarcity of large-scale, high-quality data for training and evaluating models in Olympiad-level math reasoning. In this paper, we utilize the posts from the AoPS forum to create a large-scale training and a contamination-resistant evaluation set. Our pipeline is designed to run automatically, enabling us to build and maintain evolving train/evaluation datasets. This automated approach is crucial, as it allows for continuously updating the datasets, ensuring they are less likely to suffer from contamination, even as existing data potentially becomes compromised over time. In summary, our key contributions are as follows:
We build a pipeline to extract questions and solutions from raw AoPS forum data, constructing the AoPS-Instruct, a novel large-scale dataset with K Olympiad-level math QA pairs.
Using the most recent QA pairs, We build an automatic pipeline that introduces LiveAoPSBench, a contamination-resistant evaluation set for assessing the math reasoning capabilities of LLMs. LiveAoPSBench stays regularly updated.
Our experiments on LiveAoPSBench show a declining performance trend over time for various LLMs, indicating potential data contamination, and stressing the need for up-to-date evaluation data.
Fine-tuning various LLMs on AoPS-Instruct lead to improved performance on standard benchmarks such as OlympiadBench, Omni-Math, and our LiveAoPSBench dataset, verifying the effectiveness of our dataset in enhancing math reasoning capabilities of LLMs.
2 Related Work
In this section, we provide an overview of the existing mathematical datasets used for evaluation and training purposes. Additionally, we review the latest methods and LLMs for enhancing and evaluating these math datasets.
| Model | 2023 | 2024 | Drop | |
| (K) | (K) | (%) | ||
| 7B-Qwen2.5-Math-Ins | 34.80 | 33.40 | 4.02 | |
| 72B-Qwen2-Math-Ins | 37.84 | 36.15 | 4.45 | |
| 72B-Qwen2.5-Math-Ins | 42.36 | 40.45 | 4.51 | |
| 72B-NuminaMath-CoT | 25.59 | 24.14 | 5.68 | |
| 20B-Internlm2-Math-Plus | 17.78 | 16.03 | 9.83 | |
| 16B-DeepSeekCoderV2-Ins | 22.08 | 19.80 | 10.31 | |
| 7B-Qwen2-Math-Ins | 33.26 | 29.32 | 11.85 | |
| 7B-NuminaMath-CoT | 16.88 | 14.76 | 12.55 | |
| 7B-DeepSeek-Math-RL | 14.35 | 12.44 | 13.35 | |
| 7B-Mathstral-v0.1 | 15.25 | 13.00 | 14.76 | |
| 7B-Internlm2-Math-Plus | 16.26 | 13.64 | 16.16 | |
| 27B-Gemma2-it | 12.78 | 11.59 | 9.30 | |
| 70B-Llama-3.1-Ins | 22.02 | 19.34 | 12.16 | |
| 8B-Llama3.1-Ins | 13.01 | 10.85 | 16.55 | |
| 3B-Llama3.2-Ins | 12.67 | 10.32 | 18.51 | |
| 9B-Gemma2-it | 11.63 | 9.30 | 20.01 | |
| 1B-Llama-3.2-Ins | 6.32 | 4.83 | 23.62 |
Evaluation Datasets for Math. The evaluation of the mathematical capabilities of LLMs has traditionally relied on well-established and widely-used datasets such as GSM8K and MATH (Cobbe et al., 2021; Hendrycks et al., 2021b), which have served as benchmarks for several years. These datasets typically contain math problems ranging from middle-school to high-school level, providing broad coverage across various problem categories. However, they present two significant limitations: 1) being older, their test sets are more susceptible to contamination from current training data of LLMs (Yang et al., 2023), and 2) they have reached a level of saturation, with state-of-the-art (SOTA) models achieving over 90% accuracy (Yang et al., 2024b). To address these shortcomings, Zhang et al. (2023b) introduced the Gaokao dataset, which includes more challenging high school-level problems from the Chinese college entrance exam. In addition, newer datasets such as OlympiadBench (He et al., 2024), AMC23 (AOPS, 2023), AIME24 (AOPS, 2024), and Omni-Math (Gao et al., 2024) represent higher levels of difficulty, collecting from more recent high school competition problems. While these datasets temporarily mitigate the risk of data contamination, they remain susceptible to this issue as LLMs continue to evolve, particularly with fine-tuning on newer data. To address this, we introduce LiveAoPSBench, which utilizes the most recent posts from the AoPS forum and applies substring-matching techniques to exclude any previously used problems from the new posts. More importantly, our pipeline is fully automated, allowing the evaluation set to evolve with forum posts, thereby significantly decreasing the likelihood of contamination.
Training Datasets for Math. Training datasets for mathematical reasoning can be categorized into two types: pretraining and supervised fine-tuning (SFT) datasets. First, pretraining datasets consist of large-scale math data, e.g., billions of tokens used during the pretraining phase of LLMs. Notable examples include Open-Web-Math (Paster et al., 2024) and Minerva (Lewkowycz et al., 2022), which contain 38.5B and 14.7B tokens of math data, respectively. Second, SFT datasets focus on high-quality question-answer pairs. Examples include Open-Math-Instruct (Toshniwal et al., 2024), Orca-math (Mitra et al., 2024), MetaMath (Yu et al., 2024), DART-Math (Tong et al., 2024), ScaleQuest (Ding et al., 2024), (Shah et al., 2024), and the training sets of widely used benchmarks such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b). However, these datasets are generally limited to grade-school and intermediate high-school level mathematics and do not target more advanced topics like Olympiad-level math. One of the most closely related datasets to ours is Numina (Li et al., 2024), which combines popular SFT datasets like Orca-math, MATH, and GSM8K, along with approximately 190K new Olympiad-level QA pairs. Concurrently, Yue et al. (2024) introduced a large-scale instruction fine-tuning dataset for math and science, which has also shown improvements in mathematical reasoning. Table 1 presents a detailed comparison of our dataset with these related datasets.
| Dataset | Dataset Size | Time | Olympiad | Solution | |
| Train | Eval | Stamp | Level | Rewrite | |
| Numina | K⋆ | K | ✗ | ✓ | GPT4-o |
| OpenMathInstruct | M | ✗ | ✗ | Mixtral† | |
| OlympiadBench | K | ✗ | ✓ | Human | |
| DART-Math | K | K | ✗ | ✗ | DeepSeekMath† |
| GSM8K | K | K | ✗ | ✗ | Human |
| MATH | K | K | ✗ | ✗ | Human |
| Orca-Math | K | - | ✗ | ✗ | GPT-4† |
| MAmmoTH2 | M | - | ✗ | ✗ | Mixtral |
| AoPS (Ours) | K | K | ✓ | ✓ | Qwen 2.5 |
Contamination-Resistant Evaluation. Benchmarks that are publicly accessible are prone to be contaminated due to the potential inadvertent data overlap during training. The typical decontamination method involves using exact substring (e.g., -gram) matching to detect overlaps with the target evaluation sets (Zhuo et al., 2024). However, this approach fails to catch rephrased examples and can not eliminate all overlaps with the test set (Yang et al., 2023). While alternative LLM-based methods for decontamination have been proposed, they often lack guarantees and may result in high false-positive rates (Yang et al., 2023). A reliable way to mitigate contamination is to select data that appeared after LLMs were trained, known as the knowledge cut-off. In the code generation domain, LiveCodeBench (Jain et al., 2024) addresses this issue by categorizing data based on timestamps, setting a cutoff date, and designating data beyond this point as unseen. We adopt a similar strategy in the math domain, partitioning the dataset by timestamps and enabling users to select data based on specific dates. Although this approach may not fully eliminate rephrased existing questions, it ensures that evaluation data remains unseen and less contaminated, providing a more accurate and fair assessment of LLMs.
Math-Specific Models. Several specialized models have been developed to improve the mathematical reasoning capabilities of LLMs (Shao et al., 2024; Mistral, 2024; Li et al., 2024; Yang et al., 2024b; Azerbayev et al., 2024). These models are typically initialized from pretrained general-purpose models, trained on large math datasets, followed by math-specific SFT, and then refined through reinforcement learning with human feedback (RLHF). In this paper, we fine-tune both general and math-specific models to demonstrate that AoPS-Instruct brings consistent improvements.
3 AoPS Dataset
In this section, we first describe the process of extracting and cleaning QA pairs from the AoPS forum to construct our training set. Then we explain how to utilize the latest forum data to create a reliable, contamination-resistant evaluation dataset for assessing model performance.
3.1 Math Instruction Fine-tuning Dataset: AoPS-Instruct
We now describe the five steps of our automated pipeline for constructing the instruction fine-tuning dataset AoPS-Instruct.
Step 0: Raw Forum Discussion Collection. We begin by collecting raw discussions from the forum website, where each discussion is called a “topic”. In these topics, the author presents math problems (e.g., competition-level problems) or general questions, such as seeking advice or resources. Our raw dataset consists of topics. Topics posted up until December 2023 are used as the training set, while those posted between January and August 2024 are reserved as the evaluation dataset.
Step 1: Math Question Detection. We then filter out irrelevant topics, specifically those not containing a mathematical question. To achieve this, we use Qwen 2.5 14B (Yang et al., 2024a) to decide the relevance of each topic. The first post of each topic determines whether the topic is a mathematical question or not, so we manually design a few-shot prompt, provide the first post of the topic to the model, and prompt the model to output if the topic is a math question or not. This step reduces the dataset to topics with math questions after pruning irrelevant ones.
Step 2: Question-Answer Extraction. After filtering, we extract the math question from the first post of each topic and identify potential solutions provided in subsequent posts. Since this task requires understanding the entire conversation and determining which responses contain valid solutions, we employ the 70B variant of Llama 3.1 for this step, enabling the detection of both the question and all relevant answers from the discussion.
Step 3: Solution rewriting. Math solutions generated by users on the AOPS forum are often concise, omitting details assumed to be common knowledge among the target audience. For instance, a user might write without explicitly mentioning the application of the AM-GM inequality to . While such brevity is typical for expert-level discussions, LLMs trained on these succinct solutions often struggle to maintain their chain-of-thought reasoning capabilities.
Our experiments show that fine-tuning a model on these concise solutions significantly degrades its performance on standard benchmarks (see Section 4.5 and Figure 5(b)). To address this issue, we utilize the Qwen 2.5 72B model (Yang et al., 2024b) to rewrite all solutions into detailed, step-by-step explanations. This approach aligns with similar techniques used in prior work, such as the Numina project (Li et al., 2024), which also employed solution rewriting to improve response quality. An example of a rewritten solution is provided in Figure 3, and the overall dataset curation process is illustrated in Figure 2.
Step 4: Data Decontamination. After processing all the QA pairs, we apply data decontamination to remove any overlap with the test sets of commonly used math benchmarks. Following the approach used in DeepSeekMath (Shao et al., 2024), we employ a 10-gram exact match decontamination (Zhuo et al., 2024) method to ensure that our dataset remains distinct from those benchmarks.
3.2 Contamination-Resistant Evaluation: LiveAoPSBench
Math LLMs are trained on large instructional corpora. A common issue with current evaluation sets is the risk of contamination, where test samples may inadvertently overlap with training data. To create contamination-resistant benchmarks, we constructed our evaluation set by sorting the raw data based on the initial posting timestamp and including only the most recent entries. Our evaluation set, denoted as LiveAoPSBench, is sourced from the AoPS forum, with posts strictly between January 2023 and September 2024. We utilize the same pre-processing pipeline, depicted in Figure 2, to extract QA pairs and have the raw solutions rewritten for consistency.
Filtering. The correctness of the solution is typically verified by comparing the final answer to the human-annotated answer. Note that human-annotated answers may still contain errors, as we do not perform formal proofs or verification. When constructing an evaluation set, it is essential that each question has a concrete and definite answer, which is enclosed as format for ease of parsing, as illustrated in Figure 3. We start by applying a series of heuristic filters to exclude proof-based questions and extract only those with explicit, boxed answers. To ensure that our test set does not contain problems included in widely used training sets, we use an stricter 8-gram matching filter—stricter compared to the 10-gram filter used for training set decontamination. This helps eliminate any potential overlap with common training corpora (Hendrycks et al., 2021b; Cobbe et al., 2021; Mitra et al., 2024).
Cross-Check by LLMs. A key challenge in building a fair evaluation set is ensuring the accuracy and validity of QA pairs. To automate this process, we employed two different models—Llama3.1-70B-Ins (Dubey et al., 2024) and Qwen2.5-72B-Ins (Yang et al., 2024a) to perform the rewriting step twice for each question. Consequently, for each question , we obtain a triplet: . If a boxed answer is detected in , it is automatically accepted as a candidate answer for the question. Following this, we performed a cross-check between and , removing all cases with inconsistent answers. This was done through string matching for text and value matching for numbers, while a SymPy-based (Meurer et al., 2017) symbolic equivalence program was used for SymPy-parsable expressions. The final answers are obtained by deduplicating the candidate answers. Through this process, we constructed LiveAoPSBench, which contains 3863 222The version of LiveAoPSBench we present in this paper corresponds to LiveAoPSBench-0824 which contains problems posted between 01-2024 to 08-2024. examples, all of which are from posts in the year 2024. Further details can be found in Appendix A.
Quality Verification. We assess the quality of our dataset by having a group of 10 graduate students annotate a randomly selected 10% subset (386 cases) from our evaluation set. Each human annotator verifies whether the final answer is correct based on the raw post, with each question annotated by two different individuals. We report the percentage of cases marked as correct by the human annotators to measure the correlation between human judgment and our method. Additionally, since Olympiad-level questions can be challenging even for humans, we also report the inter-annotator agreement to evaluate consistency between different groups of human annotators. More details can be found in Section 4.5.
Evolving Evaluation with Up-to-date Data. Since our pipeline does not require human annotators, we are able to continuously update our LiveAoPSBench in an automated manner. This makes our benchmark an up-to-date and timestamped evaluation set that is resistant to contamination, thereby providing a more reliable mathematical evaluation resource for the research community.
4 Experiments
4.1 Dataset Statistics
We provide a better overview of the AoPS dataset in Figure 4. As shown in Figure 4(a), more than of the questions have only one answer, while around and have two and three answers, respectively. Figure 4(b) shows the number of posts across each year, with a cut-off of August 2024. We observe that each year at least K mathematical questions are posted to the forum. This translates to more than monthly questions, which shows the potential of the AoPS forum to be used as training, and especially evaluation set. Figure 4(c) shows a breakdown of the types of questions in our dataset. Proof questions and numerical questions with about and constitute the majority of the questions in our dataset.
Finally, Figure 4(d) shows the pairwise overlap between each pair of popular supervised fine-tuning datasets using substring matching between the two datasets of each pair. Among the two Olympiad-level datasets (i.e., ours and Numina), our dataset has the least overlap with common datasets (with less than overlap), which shows the number of new data points.
4.2 Evaluation Set with Timestamps Mitigate Contamination
| Time | 23/01-04 | 23/05-08 | 23/09-12 | 24/01-04 | 24/05-08 |
| Overlap (%) | 13.24 | 11.65 | 12.82 | 9.92 | 6.88 |
| Count | 229 | 208 | 218 | 226 | 109 |
We analyze the relationship between timestamps and contamination using our evaluation set, tested against the Numnia Training Set (released July 2024). As shown in Table 2, we report contamination levels based on 10-gram overlap rates across different timestamps. We observe a consistent decrease in contamination rate as the test data timestamps become more recent.
4.3 Evaluating Open-Sourced Models
We evaluate the models’ performance as a function of time window. As shown in Fig 1, we find that all the models experience a performance drop when evaluating 2024 questions compared to questions in 2023. This decline suggests that performance on earlier examples may not accurately reflect the true capabilities of LLMs, as the initial results could be inflated by inadvertent data overlap.
4.4 Instruction Fine-Tuning
| Model | SFT Dataset | AoPS24 | Math | Olympiad Bench | Omni Math |
| No SFT | 11.7 | 47.1 | 14.5 | 12.3 | |
| Deepseek-Math | Numina | 16.3 | 55.5 | 22.7 | 17.0 |
| 7b-Ins | AoPS-Ins | 19.0 | 58.8 | 24.3 | 17.8 |
| Numina+AoPS-Ins | 19.7 | 58.8 | 25.6 | 18.0 | |
| No SFT | 15.40 | 56.30 | 21.20 | 15.90 | |
| Mathstral | Numina | 16.60 | 54.60 | 23.40 | 17.10 |
| 7B | AoPS-Ins | 23.60 | 60.80 | 27.10 | 19.90 |
| Numina+AoPS-Ins | 24.90 | 59.60 | 29.60 | 21.10 | |
| No SFT | 12.0 | 47.4 | 16.1 | 12.9 | |
| Llama-3.2 | Numina | 12.9 | 49.5 | 19.3 | 14.4 |
| 3B-Ins | AoPS-Ins | 16.7 | 54.6 | 19.6 | 16.4 |
| Numina+AoPS-Ins | 17.4 | 55.6 | 22.8 | 17.2 | |
| No SFT | 5.30 | 28.80 | 4.70 | 7.00 | |
| Llama-3.2 | Numina | 8.00 | 32.70 | 6.40 | 9.70 |
| 1B-Ins | AoPS-Ins | 10.00 | 34.70 | 11.10 | 11.00 |
| Numina+AoPS-Ins | 11.20 | 36.60 | 12.00 | 11.70 |
We show that the collected training dataset is effective at improving the math reasoning capabilities of LLMs. To this end, we choose representative LLMs and fine-tune them on our dataset combined with the Numina (Li et al., 2024) dataset, and show that such a combination provides superior performance compared to training on either of the datasets alone.
We use the following set of diverse models for fine-tuning evaluation: (1) Mathstral-7B (Mistral, 2024): a math-specialized model derived from Mistral-7B (Jiang et al., 2023), (2) DeepSeekMath-7B (Shao et al., 2024): a math-specialized model based on the DeepSeek family, and (3) Llama 3.2 3B (Dubey et al., 2024) and (4) Llama 3.2 1B (Dubey et al., 2024), two recent general state-of-the-art models. For each QA pair, only the question is used as the instruction, with the rewritten solution serving as the response, formatted within the model’s respective chat template. For instance, with Mathstral, we use the prompt: <s>[INST] question [/INST]solution for instruction tuning.
Consistent with prior work, we train each model for three epochs (Shao et al., 2024; Yang et al., 2024b), as we observe additional epochs provide no further benefit (see Figure 10 in the Appendix for ablation studies on the number of epochs). We explore three data mixtures for fine-tuning: (1) AoPS alone, (2) Numina alone, and (3) AoPS + Numina. For each dataset, we use the full decontaminated dataset for fine-tuning (i.e., Numina with K QA pairs, and AoPS-Instruct with K QA pairs). After fine-tuning each model, we evaluate the performance of each model on the following standard competition-level benchmarks: (1) OlympiadBench (He et al., 2024), which is an Olympiad-level evaluation dataset. Following prior literature (Yang et al., 2024a), we take only the math questions which have final answers and do not contain images or figures. This leaves us with samples from this dataset (2) Omni-MATH (Gao et al., 2024), which is a collection of problems from various mathematical olympiad competitions. (3) LiveAoPSBench set for the year 2024. The results are shown in Table 3. As shown by the table, fine-tuning with our dataset consistently boost the performance.
4.5 Ablation Studies
Evaluation Quality Assessment. We assess the quality of our evaluation set in two ways: by measuring its correlation with a well-established dataset and through manual evaluation over a subset of the data. First, He et al. (2024) compiled an Olympiad-level math evaluation set using manual assessment, which we leverage in our context to verify the quality of our method through the correlation between accuracies. Figure 5(a), demonstrates that the evaluation on LiveAoPSBench is highly correlated with carefully established benchmarks such as OlympiadBench. This demonstrates that our automatically generated benchmark aligns closely with the quality of those created through extensive human effort. Next, we subsample 10% of our evaluation set and ask human annotators to verify the correctness of the final parsed answers by referring to the original post. Annotators are given three options: yes, no, and no-answer. “Yes” and “no” indicate whether the answer is deemed correct, while “no-answer” is selected when a concrete answer is not appropriate (e.g., abstract concept questions answered with concrete examples). As a result, we found that 92% of the annotations were marked as correct, while 5% were incorrect and 3% fell under the no-answer category.
Rewritting’s effect on performance. We also ablate the effect of solution rewriting, which is an important part of our pipeline. As shown in Figure 5(b), rewriting solutions into a step-by-step format substantially improves the test accuracy across all benchmarks. The Qwen-2.5 72B based rewriting performs favorably against Llama-3.1 70b based rewriting on competition-level math benchmarks, while being slightly worse on easier grade-school math. Overall, we found Qwen to be a stronger model, providing a higher amount of details and being less verbose compared to Llama in rewriting solutions (see Figure 18 in the Appendix for a qualitative example). This suggests that rewriting solutions with stronger models can significantly improve performance on benchmarks.
5 Limitations
Absence of Visual Content. Our dataset currently focuses on text-only problems, which may limit its effectiveness in certain areas, particularly geometry. Many geometry problems rely heavily on diagrams to fully convey the problem statement. Incorporating relevant images and figures could significantly enhance the dataset’s comprehensiveness and applicability, especially in visually-dependent mathematical domains.
Evaluation of Proof-based Questions. Our evaluation dataset focuses on QA pairs with clear, final answers, which is well-suited to a broad range of Olympiad-level problems. However, a significant portion of such types of problems involve more complex proof-based questions that require detailed logical reasoning and multiple steps. While we incorporate proof-based questions in our instruction-tuning pipeline, the current evaluation pipeline lacks the ability to evaluate such questions effectively.
Quality Variability in Community-Generated Content. The community-driven content from the AoPS forum provides a rich source of high-quality data. Nevertheless, as with any community-generated content, the quality of answers and solutions can vary. While our filtering and refinement processes have successfully mitigated much of this noise, incorporating more advanced techniques in future iterations could result in better consistency and precision.
6 Conclusion and Future Work
In conclusion, this paper introduces the AoPS-Instruct dataset and LiveAoPSBench, leveraging community-driven content from the Art of Problem-Solving forum to address the challenges of limited training data and unreliable evaluation for LLMs solving Olympiad-level math problems. By developing a scalable and automated pipeline for extracting and refining question-answer pairs, this work presents a dataset containing over QA pairs, along with an up-to-date, contamination-resistant evaluation benchmark. Our experiments demonstrate significant performance improvements across multiple standard benchmarks for models fine-tuned on the AoPS-Instruct, highlighting enhanced mathematical reasoning capabilities. Furthermore, the observed performance decline of various LLMs on LiveAoPSBench underscores the importance of continuously updating evaluation sets to mitigate the risks of data contamination.
For future work, there are several promising directions to explore. First, while this paper focuses on the AoPS forum, the pipeline developed is not limited to this domain. It is generalizable and can be applied to other online forums or different subject areas, enabling the creation of high-quality datasets for various fields, such as physics, computer science, or even non-technical disciplines. Expanding this pipeline to other knowledge-intensive communities could further improve the training and evaluation of LLM across disciplines. Additionally, the quality of the dataset can be significantly improved by incorporating more advanced LLMs into the pipeline. Leveraging state-of-the-art models for question extraction, answer detection, and solution rewriting would result in more accurate and detailed data, ultimately enhancing the effectiveness of the fine-tuned models. Moreover, the problems is AoPS-Instruct could be used as a diverse set to distill larger models into smaller models. Lastly, the current pipeline focuses on question-answer pairs with clear final answers, but a significant portion of Olympiad-level problems involves proof-based questions that require a deeper evaluation of logical reasoning, argument structure, and intermediate steps. Future work could include adapting the pipeline to accommodate these proof-based problems, potentially using another advanced LLM as a judge (Li et al., 2023), or incorporating formalization methods to better assess these complex solutions.
Impact Statement
This paper aims to advance the field of mathematical reasoning. It is important to note that AoPS-Instruct and LiveAoPSBench contain no personal information. Additionally, we have not identified any societal consequences that require special attention in this context.
Acknowledgments
This work was funded, in part, by the NSERC DG Grant (No. RGPIN-2022-04636), the Vector Institute for AI, Canada CIFAR AI Chair, NSERC Canada Research Chair (CRC), NSERC Discovery Grants, a Google Gift Fund, and the Government of Canada’s New Frontiers in Research Fund NFRFE-2023-00936. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through the Digital Research Alliance of Canada alliance.can.ca, and companies sponsoring the Vector Institute www.vectorinstitute.ai/#partners, and Advanced Research Computing at the University of British Columbia. Additional hardware support was provided by John R. Evans Leaders Fund CFI grant. Sadegh Mahdavi and Muchen Li are supported by UBC Four Year Doctoral Fellowships.
Appendix A More details on LiveAoPSBench
A.1 Evaluation pipeline statistics
To begin with, we have 14158 QA pairs with time stamps between Jan-2024 and Aug-2024. Decontamination with 8-gram matching is performed against Math and GSM8K training set (Hendrycks et al., 2021b; Cobbe et al., 2021), which removes 664 Q-A pairs. After removing proof questions and non-boxed solutions, we are left with 7173 Q-A pairs over 5416 unique questions. Lastly, The LLM cross-check filters out 1553 questions with inconsistent solutions and the resulting LiveAoPSBench contains 3863 questions. We apply the same pipeline described in Sec 3.2 to data with a time stamp between Jan-2023 and Dec-2023 and get 5216 questions for the 2023 split result.
A.2 Human Annotation
As shown in Figure 6, we develop a simple web interface for human annotators to verify the answers extracted by our LLMs. Annotators compare the “Voted Answer”, “Original Answers” and all posts in the original topic page identified by LLMs to verify if the “Voted Answer” matches the original posts’ answers. The verification process provides four results: Positive (“Yes”), negative (“No/No Answer”), and neutral (“Not sure”). The “Not sure” option is provided since verifying the answer sometimes requires a certain mathematical foundation and a significant amount of reading time. We also show highlight two examples of disagreement in Figure 7.
A.3 Derivation of Difficulty Levels
The difficulty levels in this dataset do not reflect the exact difficulty of the problems but rather approximate the general education background of the problem, e.g., this is a “High School” level problem. However, a challenging high school problem may be more complex than an easy college-level problem. The classification is derived from the problem tag in the AOPS forum, where the categories correspond to “Middle School”, “High School”, “College”, and “High School Olympiads”. In addition, some problems originate from special forums, which do not fit into the above categories and are classified as “Others” in our dataset.
Example 1
Example 2
Appendix B Detailed Evaluation Results on LiveAoPSBench
B.1 LiveAoPSBench-2024
We update the LiveAoPSBench with the complete 2024 data. This resulted in LiveAoPSBench-2024 with 5328 questions. In figure 8 we provide the latest result with the state-of-the-art reasoning models like deepseek-r1(DeepSeek-AI, 2025) and QwQ(Team, 2024).
B.2 Evaluating open-sourced LLMs
We have selected several mainstream open-source general LLMs and math-specific LLMs that demonstrate high performance on the previous math evaluation datasets. For math-specific LLMs, we choose DeepSeek-Math-7b-rl (Shao et al., 2024), Mathstral-7B-v0.1 (Mistral, 2024), 7b and 20b versions of Internlm2-Math-plus (Ying et al., 2024), 7B and 72B versions of NuminaMath-CoT (Li et al., 2024), 1.5B,7B,72B version of Qwen2-Math-Instruct (Yang et al., 2024a) and Qwen2.5-Math-Instruct (Yang et al., 2024b) as the representative of the math specific LLMs. Additionally, we include DeepSeek-Coder-V2-Lite-Instruct (Zhu et al., 2024), which is a code specialist model trained on both math and code corpus. For general purpose LLMs, We report performance on 1B, 3B and 8B versions of the Llama3 family models (Dubey et al., 2024) as well as 9B and 27B versions of Gemma-2-Instruct (Team et al., 2024) model.
B.3 Detailed Results
The accuracy comparison for these mainstream open source LLMs are shown in Tables 4, 5, 6 split by Month, Difficulty and Answer Type. The Month tables separately include evaluation results for 2023 and 2024. For the Difficulty and Answer Type tables, we use only the most recent evaluation results from 2024. Notably, the difficulty labels represent the general educational background of the problems rather than their exact difficulty. Over half of the problems originate from educational backgrounds associated with High School or High School Olympiads, and only around 7% are from Middle School, indicating our dataset’s focus is more on the complex problems. Similarly, in the Answer Type Table, more than half of the problems are categorized as numeric-int.
| Model | 2023 | Jan | Feb | Mar | Apr | May | Jun | Jul | Aug | Sep | Oct | Nov | Dec |
| Count | 5216 | 483 | 388 | 444 | 415 | 472 | 412 | 396 | 505 | 381 | 409 | 404 | 507 |
| DeepSeek-Coder-V2-Lite-it | 22.45 | 23.40 | 19.33 | 22.97 | 22.17 | 22.67 | 21.12 | 23.99 | 25.74 | 22.83 | 23.72 | 22.28 | 18.93 |
| DeepSeek-Math-7b-rl | 15.38 | 18.22 | 15.46 | 17.57 | 13.49 | 14.41 | 12.86 | 15.66 | 14.65 | 15.75 | 15.89 | 17.08 | 13.61 |
| Internlm2-Math-plus-20b | 18.23 | 18.84 | 16.75 | 17.79 | 18.31 | 20.55 | 17.72 | 16.41 | 18.42 | 17.06 | 21.27 | 18.81 | 16.57 |
| Internlm2-Math-plus-7b | 17.10 | 17.81 | 16.49 | 19.82 | 17.11 | 18.01 | 15.05 | 19.95 | 16.24 | 13.91 | 18.58 | 17.08 | 15.19 |
| Mathstral-7B-v0.1 | 15.91 | 17.18 | 14.95 | 15.54 | 14.22 | 16.31 | 16.75 | 16.67 | 18.22 | 13.91 | 14.18 | 18.32 | 14.20 |
| NuminaMath-72B-CoT | 26.15 | 29.40 | 25.77 | 23.87 | 25.30 | 29.87 | 22.57 | 25.51 | 25.54 | 24.41 | 27.87 | 25.50 | 27.02 |
| NuminaMath-7B-CoT | 17.29 | 18.43 | 14.95 | 18.24 | 15.90 | 19.70 | 15.53 | 17.68 | 17.43 | 15.22 | 18.83 | 19.80 | 15.38 |
| Qwen2-Math-1.5B-it | 29.06 | 31.47 | 26.55 | 31.31 | 31.08 | 28.81 | 26.21 | 26.77 | 29.90 | 27.82 | 29.58 | 32.43 | 26.43 |
| Qwen2-Math-72B-it | 37.96 | 41.41 | 36.34 | 37.39 | 38.31 | 38.35 | 37.14 | 37.37 | 38.61 | 34.91 | 40.10 | 39.60 | 35.50 |
| Qwen2-Math-7B-it | 33.07 | 33.13 | 35.31 | 33.33 | 31.33 | 36.86 | 28.64 | 35.10 | 33.27 | 30.45 | 32.52 | 36.39 | 30.57 |
| Qwen2.5-Math-1.5B-it | 34.72 | 36.02 | 32.47 | 35.81 | 34.94 | 37.50 | 32.04 | 31.31 | 36.04 | 33.86 | 36.67 | 35.15 | 33.73 |
| Qwen2.5-Math-72B-it | 42.04 | 44.31 | 38.40 | 41.22 | 41.93 | 45.55 | 44.17 | 40.15 | 42.18 | 35.43 | 46.21 | 43.32 | 40.43 |
| Qwen2.5-Math-7B-it | 34.87 | 35.82 | 31.70 | 36.26 | 35.42 | 37.71 | 29.61 | 33.59 | 37.43 | 32.28 | 34.72 | 37.87 | 34.52 |
| Llama-3.2-1B-it | 6.75 | 6.83 | 5.93 | 4.73 | 5.54 | 9.53 | 5.10 | 8.08 | 6.73 | 8.66 | 7.09 | 7.92 | 5.13 |
| Llama-3.2-3B-it | 13.77 | 14.49 | 13.40 | 12.16 | 13.98 | 14.19 | 13.11 | 16.92 | 12.08 | 14.44 | 15.16 | 13.61 | 12.43 |
| Llama-3.1-8B-it | 14.03 | 15.53 | 13.66 | 14.86 | 13.73 | 13.56 | 11.65 | 15.15 | 16.04 | 12.86 | 14.91 | 14.60 | 11.64 |
| Gemma-2-27b-it | 13.80 | 11.39 | 13.92 | 15.77 | 13.01 | 15.04 | 10.68 | 14.90 | 15.64 | 12.86 | 14.91 | 14.36 | 13.02 |
| Gemma-2-9b-it | 12.42 | 10.97 | 10.05 | 13.51 | 12.53 | 13.35 | 11.17 | 13.38 | 15.05 | 10.76 | 13.45 | 13.61 | 10.85 |
| Model | 2024 | Jan | Feb | Mar | Apr | May | Jun | Jul | Aug |
| Count | 3863 | 634 | 527 | 614 | 503 | 511 | 380 | 363 | 331 |
| DeepSeek-Coder-V2-Lite-it | 20.86 | 18.14 | 22.77 | 25.41 | 22.47 | 18.00 | 17.89 | 21.21 | 19.64 |
| DeepSeek-Math-7b-rl | 13.64 | 14.04 | 17.08 | 15.31 | 15.71 | 10.37 | 10.53 | 13.50 | 9.97 |
| Internlm2-Math-plus-20b | 16.93 | 14.51 | 22.01 | 17.59 | 17.50 | 14.48 | 14.74 | 18.73 | 15.71 |
| Internlm2-Math-plus-7b | 14.81 | 11.83 | 18.60 | 17.59 | 16.50 | 11.15 | 17.11 | 13.77 | 10.88 |
| Mathstral-7B-v0.1 | 14.29 | 11.04 | 17.65 | 17.43 | 16.50 | 10.96 | 12.11 | 12.95 | 15.11 |
| NuminaMath-72B-CoT | 24.95 | 22.08 | 28.65 | 28.66 | 23.46 | 23.48 | 22.37 | 27.55 | 22.36 |
| NuminaMath-7B-CoT | 16.13 | 13.56 | 20.87 | 17.10 | 19.68 | 13.89 | 13.42 | 16.25 | 12.69 |
| Qwen2-Math-1.5B-it | 26.84 | 25.24 | 29.79 | 30.46 | 27.24 | 22.90 | 25.79 | 28.37 | 23.56 |
| Qwen2-Math-72B-it | 36.68 | 34.38 | 37.57 | 41.69 | 36.98 | 36.59 | 35.79 | 36.91 | 30.82 |
| Qwen2-Math-7B-it | 30.05 | 28.08 | 31.69 | 34.20 | 31.41 | 27.98 | 29.74 | 30.85 | 24.17 |
| Qwen2.5-Math-1.5B-it | 32.57 | 28.71 | 33.59 | 39.25 | 31.21 | 31.70 | 30.00 | 34.99 | 29.61 |
| Qwen2.5-Math-72B-it | 40.56 | 40.54 | 41.56 | 45.11 | 40.16 | 37.96 | 40.53 | 42.42 | 33.23 |
| Qwen2.5-Math-7B-it | 34.04 | 32.18 | 35.48 | 41.37 | 32.41 | 32.09 | 32.11 | 34.99 | 28.40 |
| Llama-3.2-1B-it | 5.80 | 3.63 | 7.40 | 6.84 | 6.16 | 4.70 | 5.53 | 7.16 | 5.44 |
| Llama-3.2-3B-it | 11.75 | 8.68 | 13.85 | 12.87 | 14.12 | 10.18 | 13.42 | 9.92 | 11.18 |
| Llama-3.1-8B-it | 12.71 | 10.57 | 16.13 | 14.17 | 11.33 | 12.33 | 12.63 | 12.67 | 11.48 |
| Gemma-2-27b-it | 12.87 | 10.57 | 16.13 | 13.03 | 13.52 | 11.94 | 11.58 | 14.05 | 12.39 |
| Gemma-2-9b-it | 10.67 | 8.68 | 11.95 | 13.36 | 10.93 | 8.02 | 8.16 | 14.05 | 10.27 |
| Model | Overall | Middle School | High School | College | High School | Others |
| Olympiads | ||||||
| Count | 3863 | 286 | 1349 | 314 | 889 | 1025 |
| DeepSeek-Coder-V2-Lite-it | 20.86 | 24.48 | 19.79 | 22.93 | 17.21 | 23.80 |
| DeepSeek-Math-7b-rl | 13.64 | 22.73 | 12.23 | 14.97 | 8.55 | 16.98 |
| Internlm2-Math-plus-20b | 16.93 | 24.83 | 15.64 | 19.11 | 12.71 | 19.41 |
| Internlm2-Math-plus-7b | 14.81 | 20.63 | 13.94 | 17.20 | 9.67 | 18.05 |
| Mathstral-7B-v0.1 | 14.29 | 19.23 | 12.90 | 15.61 | 10.24 | 17.85 |
| NuminaMath-72B-CoT | 24.95 | 33.57 | 25.28 | 26.75 | 18.45 | 27.22 |
| NuminaMath-7B-CoT | 16.13 | 18.18 | 15.12 | 16.24 | 12.60 | 19.90 |
| Qwen2-Math-1.5B-it | 26.84 | 32.17 | 26.17 | 28.98 | 22.61 | 29.27 |
| Qwen2-Math-72B-it | 36.68 | 43.36 | 38.62 | 42.04 | 28.01 | 38.15 |
| Qwen2-Math-7B-it | 30.05 | 37.06 | 30.17 | 31.21 | 24.86 | 32.10 |
| Qwen2.5-Math-1.5B-it | 32.57 | 38.81 | 33.65 | 31.85 | 28.91 | 32.78 |
| Qwen2.5-Math-72B-it | 40.56 | 48.25 | 42.70 | 45.86 | 32.62 | 40.88 |
| Qwen2.5-Math-7B-it | 34.04 | 42.66 | 34.84 | 36.62 | 27.11 | 35.80 |
| Llama-3.2-1B-it | 5.80 | 10.14 | 4.30 | 5.41 | 3.71 | 8.49 |
| Llama-3.2-3B-it | 11.75 | 18.53 | 10.23 | 9.55 | 8.10 | 15.71 |
| Llama-3.1-8B-it | 12.71 | 17.48 | 10.60 | 15.29 | 7.54 | 17.85 |
| Gemma-2-27b-it | 12.87 | 20.63 | 11.86 | 13.38 | 7.65 | 16.39 |
| Gemma-2-9b-it | 10.67 | 15.38 | 8.30 | 12.74 | 7.65 | 14.44 |
| Model | Overall | equation | expression | list | numeric-dec | numeric-int | numeric-irr | others |
| Count | 3863 | 296 | 950 | 195 | 57 | 2114 | 176 | 75 |
| DeepSeek-Coder-V2-Lite-it | 20.86 | 18.24 | 16.00 | 20.00 | 15.79 | 24.36 | 11.93 | 21.33 |
| DeepSeek-Math-7b-rl | 13.64 | 11.15 | 10.42 | 10.77 | 8.77 | 16.18 | 6.25 | 21.33 |
| Internlm2-Math-plus-20b | 16.93 | 14.53 | 12.74 | 13.33 | 14.04 | 20.20 | 8.52 | 18.67 |
| Internlm2-Math-plus-7b | 14.81 | 11.49 | 9.26 | 12.31 | 12.28 | 18.31 | 8.52 | 22.67 |
| Mathstral-7B-v0.1 | 14.29 | 14.53 | 10.74 | 6.67 | 12.28 | 16.65 | 10.80 | 21.33 |
| NuminaMath-72B-CoT | 24.95 | 19.59 | 20.21 | 16.92 | 24.56 | 28.71 | 20.45 | 32.00 |
| NuminaMath-7B-CoT | 16.13 | 14.86 | 12.11 | 11.28 | 12.28 | 18.78 | 12.50 | 21.33 |
| Qwen2-Math-1.5B-it | 26.84 | 23.31 | 22.11 | 28.72 | 15.79 | 29.52 | 25.57 | 32.00 |
| Qwen2-Math-72B-it | 36.68 | 27.70 | 30.84 | 36.41 | 26.32 | 41.15 | 31.82 | 40.00 |
| Qwen2-Math-7B-it | 30.05 | 23.99 | 26.21 | 27.69 | 19.30 | 33.30 | 25.00 | 37.33 |
| Qwen2.5-Math-1.5B-it | 32.57 | 25.34 | 28.74 | 32.82 | 22.81 | 35.86 | 22.73 | 46.67 |
| Qwen2.5-Math-72B-it | 40.56 | 31.42 | 38.32 | 41.54 | 31.58 | 43.19 | 36.36 | 45.33 |
| Qwen2.5-Math-7B-it | 34.04 | 28.72 | 30.21 | 33.85 | 24.56 | 36.90 | 30.11 | 40.00 |
| Llama-3.2-1B-it | 5.80 | 2.70 | 3.79 | 2.56 | 5.26 | 7.66 | 3.41 | 5.33 |
| Llama-3.2-3B-it | 11.75 | 7.77 | 8.32 | 7.18 | 7.02 | 14.71 | 6.25 | 16.00 |
| Llama-3.1-8B-it | 12.71 | 4.05 | 9.58 | 10.77 | 17.54 | 15.33 | 9.66 | 21.33 |
| Gemma-2-27b-it | 12.87 | 7.77 | 10.11 | 9.74 | 14.04 | 15.28 | 9.09 | 16.00 |
| Gemma-2-9b-it | 10.67 | 7.09 | 7.89 | 8.72 | 14.04 | 12.54 | 7.95 | 16.00 |
Appendix C Training set details
C.1 Decontamination Details
We use 10-gram substring matching to decontaminate against test set for a comprehensive list of math evaluation datasets available. (Cobbe et al., 2021; Hendrycks et al., 2021b; He et al., 2024; AOPS, 2023, 2024; Zhang et al., 2023b; Lewkowycz et al., 2022; Gao et al., 2024; Miao et al., 2020; Hendrycks et al., 2021a; Koncel-Kedziorski et al., 2016; Patel et al., 2021; Zhang et al., 2023a). In Figure 9. We show the decontamination statistic for our dataset and Numina.
Appendix D SFT Experiements
D.1 Ablation with controlled computation budget
As shown in Tab 3, Numina + AoPS-Instruct performs favorably against using AoPS-Instruct or Numina alone. To show this gain is not simply achieved by doubling the computation available for fine-tuning. We control the total fine-tune budget the same for AoPS-Instruct only, numina only and AoPS-Instruct + numina. This results in approximately 6 epoch on AoPS-Instruct or Numina or 3 epoch of training on AoPS-Instruct + Numina. We show the curve of ACC on Math,LiveAoPSBench, OlympiadBench w.r.t. training steps.
D.2 Rewriting model ablation
We use Qwen 2.5 72B to rewrite the solutions, and then we fine-tune smaller models on our dataset. This may raise the question of whether the effectiveness of our dataset would be limited by the capabilities of its rewriting model. To show the effectiveness of our dataset, we use a Qwen 2.5 1.5B to rewrite the solutions and then fine-tune DeepSeek-Math 7B-instruct on the dataset. Table 7 shows the performance of the original DeepSeek-Math, the performance of Qwen2.5-1.5B, and the performance of fine-tuned DeepSeek on Qwen2.5-1B-rewritten solutions. As shown by the results, the fine-tuned version outperforms both models, which shows that our dataset can improve the reasoning capability beyond its rewriting model solution.
| Model | AIME24 | AMC23 | Olympiad Bench | Math | AoPS24 | Omni Math |
| Deekseek-Math-7b-Ins | 1/30 | 8/40 | 14.5 | 47.1 | 11.7 | 12.3 |
| Qwen2.5-1.5b-Ins | 0/30 | 9/40 | 21.3 | 55.0 | 16.7 | 16.8 |
| Deepseek-Math-7b-Ins (fine-tuned) | 1/30 | 13/40 | 22.7 | 61.0 | 19.4 | 19.2 |
Appendix E AIME and AMC results on the main models
Results of fine-tuning on AIME 2024 and AMC 2024 are present in table 8.
| Model | SFT Dataset | AIME24 | AMC23 |
| No SFT | 1/30 | 8/40 | |
| Deepseek-Math | Numina | 0/30 | 12/40 |
| 7b-Ins | AoPS-Ins | 1/30 | 15/40 |
| Numina+AoPS-Ins | 2/30 | 11/40 | |
| No SFT | 0/30 | 16/40 | |
| Mathstral | Numina | 0/30 | 15/40 |
| 7B | AoPS-Ins | 0/30 | 12/40 |
| Numina+AoPS-Ins | 0/30 | 16/40 | |
| No SFT | 2/30 | 11/40 | |
| Llama-3.2 | Numina | 1/30 | 6/40 |
| 3B-Ins | AoPS-Ins | 2/30 | 11/40 |
| Numina+AoPS-Ins | 0/30 | 12/40 | |
| No SFT | 0/30 | 5/40 | |
| Llama-3.2 | Numina | 0/30 | 6/40 |
| 1B-Ins | AoPS-Ins | 0/30 | 6/40 |
| Numina+AoPS-Ins | 0/30 | 9/40 |
| Dataset | Size | LiveAoPSBench | Olympiad Bench | Omni Math | MATH |
| DART-MATH-Hard | K | 14.4 | 21.8 | 15.4 | 52.5 |
| AoPS-Instruct | K | 19.0 | 24.3 | 17.8 | 58.8 |
Appendix F Use of AoPS as a Data Source
Concurrent to our work, both Numina (Li et al., 2024) and Omni-math (Gao et al., 2024) also use AoPS as their data source. Different from us, Numina only includes data from the contest page with 30K questions 333https://artofproblemsolving.com/community/c13_contest_collections, while we utilize all the available posts on this forum. Furthermore, Omni-math (Gao et al., 2024) includes only 4428 evaluation questions from all timestamps, while we include the most recent problems posted in 2024, as well as a large-scale training set.
Appendix G Prompts
Appendix H Dataset Examples
We provide further examples of our dataset and its rewritten solutions in Figures 14, 15, 16, 17. 18, and 19.